# Toward General-Purpose Robots via Foundation Models: A Survey and Meta-Analysis

### v2

@@ 5.2 Analysis of Current Method Evaluation
Need for Improved Generalization and Robustness:
@@
Limited Exploration in Low-Level Actions:
@@
Control Frequencies Too Slow to be Deployed on Real Robots:
@@
Lack of Uniform Benchmarks for Testing:
@@ 6.1 Remaining Challenges and Open Discussions
Although numerous strategies have been explored to address the problem of grounding, as discussed in Section 9, there are many open challenges in this area. First, grounding needs an effective medium or interface that bridges concepts and robot actions. Existing interfaces, such as those employing natural language [24, 31] and code [216, 221], are limited. While concepts can be articulated through language and code, they are not universally applicable to nuances such as dexterous body movements. Furthermore, these interfaces often depend on predefined skill libraries that are not only time-intensive to develop but also lack generalization to new environments. Using reward as an interface [226, 232, 182] may alleviate some of the generalization issues in simulations by acquiring skills dynamically. However, the time-consuming and potentially unsafe nature of training RL algorithms in the real world raises questions about the feasibility of this method, with real-world validations of its effectiveness yet to be demonstrated.
Second, we need to move from a unimodal notion of grounding, like mapping the word to meaning to a more holistic grounding of multiple sensory modalities. Approaches that rely solely on visual data [280] may capture certain physical properties such as material, transparency, and deformability. Yet, they fall short in grasping concepts like friction, which requires interactive data with proprioceptive feedback, or the scent of an object, which cannot be acquired without additional modalities such as olfaction.
@@
As we pursue deployments of real robots to work alongside humans in factories, to provide elderly care, or to assist in homes and offices, these autonomous systems (and the foundation models that power them) will require more effective measures of safety. While formal hardware and software safety checks still apply, the use of foundation models to support provable safety analysis will become an increasingly necessary direction. With the goal of deploying robots to safety-critical scenarios, prior works have considered leveraging Lyapunov-style safety index functions [323, 324, 122], in attempts to provide hard safety guarantees for nonlinear systems with complex dynamics and external disturbances (see also Section 3.5). Traditionally, the systems under consideration by the provable safety literature are of low dimension, often require careful specification of a world/dynamics model, require specifying an initial safe set and/or set-boundary distance functions, require some heuristics and training “tricks” to obtain useful safety value functions that balance conservativeness versus performance, do not naturally support multi-agent settings, and present challenges in safely updating the safety value function and growing the safe set online. Herbert et al. [324] synthesized several techniques into a framework—thereby easing computation, streamlining updates to the safe sets by one or more orders of magnitude compared to the prior art, and extending Hamilton-Jacobi Reachability analysis to 10-dimensional systems that govern quadcopter control. Chen et al. [122] combine RL with HJ Reachability analysis to learn safety value functions from high-dimensional inputs (RGB images, plus vehicle state), to trade-off a performance-oriented policy and a safety-oriented policy, within a jointly-optimized dual actor-critic framework, for simulated autonomous racing. Tian et al. [325] integrate HJ Reachability analysis in the context of multi-agent interactions in urban autonomous driving, by formulating the problem as a general-sum Stackelberg game.
@@
Regarding the architecture and optimization design for foudnation models used in robotics, we can focus on a functional approach rather than categorizing it as either modular or end-to-end differentiable. One of the goals of a robotic foundational model is to allow flexible modular components, each responsible for specific functionalities, with unified learning that leverages shared representations and general learning capabilities.
@@
one of the key open research question is how robotics foundational models should enable motion control across different physical embodiments. Robot policies deployed in homes and offices must be robust to mechanical motion failures, such as sensor malfunctions or actuator breakdowns, ensuring continued functionality in challenging environments. Furthermore, robotic systems must be designed to adapt to a variety of tools and peripherals, mirroring the human capability to interact with different instruments for specific tasks and physical tool uses. While some works [331, 332, 333] have explored learning representations for diverse tool use, these approaches are yet to be scaled up with foundation models.
@@
The foundation models surveyed in this paper mostly take a model-agnostic (model-free) approach, leveraging the strength of expansive datasets and large-scale deep learning architectures. Some exceptions have attempted to emulate model-based approaches by directly employing LLMs as dynamic models. However, these attempts are still constrained by the inherent limitations of text-only descriptions and are prone to encountering issues with hallucinations, as discussed in [271, 225]. Many researchers would argue [138] that the data-scaled learning paradigm of these foundation models is still quite different from how humanity and animals learn, which is in an extreme data- and energy-efficient manner. Achieving even close to the joint performance and efficiency of human learning ability remains intriguing. In [138], LeCun argues that one possible answer to resolving that puzzle may lie in the learning of world models, a model that predicts how the state of the world going to change as consequences of the actions taken. If we were to develop world models that can emulate the precision of the world’s representation through rigorous mathematical and physical modeling, it would bring us significantly closer to addressing and generalizing complex issues in robotics. These sophisticated and reliable world models would enable the application of established model-based methodologies, including search-based and sample-based planning, as well as trajectory optimization techniques. This approach would not only facilitate the resolution of planning and control challenges in robotics but also augment the explainability of these processes. It is posited that the pursuit of a ’foundation world model’, characterized by remarkable generalization abilities and zero-shot learning capabilities, holds the potential to be a paradigm-shifting development in the field.
@@
As demonstrated in Figure 3(c) and the Meta-analysis in Tables 2-7, existing real robot platforms utilized for deploying foundation models are predominantly limited to gripper-based, single-arm robot manipulators. The range of concepts learnable from tasks executed by these hardware systems is restricted, primarily because the simple opening and closing actions of a gripper are easily describable by language. To enable robots to achieve a level of dexterity and motor skills comparable to those of animals and humans, or to perform complex domestic tasks, it is essential for foundation models to acquire a deeper understanding of physical and household concepts. This learning necessitates a broader spectrum of information sources, such as diverse sensors (including smell, tactile, and thermal sensors), and more intricate data such as proprioception data from robot platforms with high degrees of freedom.
Current dexterous manipulators, e.g., Shadow Hand [334], are prohibitively expensive and prone to frequent breakdowns, hence they are predominantly experimented with in simulation. Moreover, tactile sensors are still limited in their application, often confined to the fingertips, as in [335], or offer only low resolution, as observed in the robot-sweater [336]. Furthermore, since the bulk of data-collection is still conducted through human demonstrations, platforms that facilitate more accurate and efficient data acquisition, such as ALOHA [337] and Leap Hands [284], are gaining popularity. Therefore, we posit that significant contributions are yet to be made—not only in terms of software innovations, but also in hardware. These advancements are crucial for providing richer data-collection and, thus, expanding the conceptual space of robotics foundation models.
@@
The robotics community needs to encourage standardized and reproducible research practices to ensure that published findings can be validated and compared by others. To enable reproducibility at scale, we need to bridge the gap between simulated environments and real-world hardware and improve the transferability of ML models. Homerobot [205] is a promising step towards enabling both simulation and hardware platforms for open vocabulary pick-and-place tasks. We need to establish standardized task definitions and affordances to handle different robot morphologies, enabling more efficient model development.
@@ 6.2 Summary
Due to the rapidly changing nature of the field, we checkpointed this version of literature review on Dec 13th 2023, and might have missed some relevant work. In addition, due to the rich body of literature and the extensiveness of this survey, there may be inaccuracies or mistakes in the paper. We welcome readers to send pull requests to our GitHub repository (inside https://robotics-fm-survey.github.io/) so we may continue to update our references, correct the mistakes and inaccuracies, as well as updating the entries of the meta studies in the paper. Please refer to the contribution guide in the GitHub repository.
@@ Acknowledgments
We would like to thank Vincent Vanhoucke for feedbacks on a draft of this survey paper. In addition, we would like to thank Kedi Xu for insightful discussions about the papers list.

### v3

@@ 5.2 Analysis of Current Method Evaluation
Need for Improved Generalization and Robustness
@@
Limited Exploration in Low-Level Actions
@@
Control Frequencies Too Slow to be Deployed on Real Robots
@@
Lack of Uniform Benchmarks for Testing
@@ 6.1 Remaining Challenges and Open Discussions
Although numerous strategies have been explored to address the problem of grounding, as discussed in Section 4.1.4, there are many open challenges in this area. First, grounding needs an effective medium or interface that bridges concepts and robot actions. Existing interfaces, such as those employing natural language [26, 33] and code [122, 121], are limited. While concepts can be articulated through language and code, they are not universally applicable to nuances such as dexterous body movements. Furthermore, these interfaces often depend on predefined skill libraries that are not only time-intensive to develop but also lack generalization to new environments. Using reward as an interface [134, 133, 78] may alleviate some of the generalization issues in simulations by acquiring skills dynamically. However, the time-consuming and potentially unsafe nature of training RL algorithms in the real world raises questions about the feasibility of this method, with real-world validations of its effectiveness yet to be demonstrated.
Second, we need to move from an unimodal notion of grounding, like mapping the word to meaning to a more holistic grounding of multiple sensory modalities. Approaches that rely solely on visual data [206] may capture certain physical properties such as material, transparency, and deformability. Yet, they fall short in grasping concepts like friction, which requires interactive data with proprioceptive feedback, or the scent of an object, which cannot be acquired without additional modalities such as olfaction.
@@
As we pursue deployments of real robots to work alongside humans in factories, to provide elderly care, or to assist in homes and offices, these autonomous systems (and the foundation models that power them) will require more effective measures of safety. While formal hardware and software safety checks still apply, the use of foundation models to support provable safety analysis will become an increasingly necessary direction. With the goal of deploying robots to safety-critical scenarios, prior works have considered leveraging Lyapunov-style safety index functions [219, 220, 88], in attempts to provide hard safety guarantees for nonlinear systems with complex dynamics and external disturbances (see also Section 3.5). Traditionally, the systems under consideration by the provable safety literature are of low dimension, often require careful specification of a world/dynamics model, require specifying an initial safe set and/or set-boundary distance functions, require some heuristics and training “tricks” to obtain useful safety value functions that balance conservativeness versus performance, do not naturally support multi-agent settings, and present challenges in safely updating the safety value function and growing the safe set online. Herbert et al. [220] synthesized several techniques into a framework—thereby easing computation, streamlining updates to the safe sets by one or more orders of magnitude compared to the prior art, and extending Hamilton-Jacobi Reachability analysis to 10-dimensional systems that govern quadcopter control. Chen et al. [88] combine RL with HJ Reachability analysis to learn safety value functions from high-dimensional inputs (RGB images, plus vehicle state), to trade off a performance-oriented policy and a safety-oriented policy, within a jointly-optimized dual actor-critic framework, for simulated autonomous racing. Tian et al. [221] integrate HJ Reachability analysis in the context of multi-agent interactions in urban autonomous driving, by formulating the problem as a general-sum Stackelberg game.
@@
Regarding the architecture and optimization design for foundation models used in robotics, we can focus on a functional approach rather than categorizing it as either modular or end-to-end differentiable. One of the goals of a robotic foundational model is to allow flexible modular components, each responsible for specific functionalities, with unified learning that leverages shared representations and general learning capabilities.
@@
One of the key open research question is how robotics foundational models should enable motion control across different physical embodiments [198]. Initial results [198] show the possibility of one model for different policies across different embodiment.
Robot policies deployed in homes and offices must be robust to mechanical motion failures, such as sensor malfunctions or actuator breakdowns, ensuring continued functionality in challenging environments. Furthermore, robotic systems must be designed to adapt to a variety of tools and peripherals, mirroring the human capability to interact with different instruments for specific tasks and physical tool uses.
@@
The foundation models surveyed in this paper mostly take a model-agnostic (model-free) approach, leveraging the strength of expansive datasets and large-scale deep learning architectures. Some exceptions have attempted to emulate model-based approaches by directly employing LLMs as dynamic models. However, these attempts are still constrained by the inherent limitations of text-only descriptions and are prone to encountering issues with hallucinations, as discussed in [195, 127]. Many researchers would argue [227] that the data-scaled learning paradigm of these foundation models is still quite different from how humanity and animals learn, which is in an extreme data- and energy-efficient manner. Achieving even close to the joint performance and efficiency of human learning ability remains intriguing. In [227], LeCun argues that one possible answer to resolving that puzzle may lie in the learning of world models, a model that predicts how the state of the world going to change as consequences of the actions taken.
If we were to develop world models that can emulate the precision of the world’s representation through rigorous mathematical and physical modeling, it would bring us significantly closer to addressing and generalizing complex issues in robotics. These sophisticated and reliable world models would enable the application of established model-based methodologies, including search-based and sample-based planning, as well as trajectory optimization techniques. This approach would not only facilitate the resolution of planning and control challenges in robotics but also augment the explainability of these processes. It is posited that the pursuit of a ’foundation world model’, characterized by remarkable generalization abilities and zero-shot learning capabilities, holds the potential to be a paradigm-shifting development in the field. And we have already seen the reality of ’foundation world model’ getting closer with action/language conditioned video generation  [228, 229]
@@
As demonstrated in Figure 4 and the Meta-analysis in Tables 2-7, existing real robot platforms utilized for deploying foundation models are predominantly limited to gripper-based, single-arm robot manipulators. The range of concepts learnable from tasks executed by these hardware systems is restricted, primarily because the simple opening and closing actions of a gripper are easily describable by language. To enable robots to achieve a level of dexterity and motor skills comparable to those of animals and humans, or to perform complex domestic tasks, it is essential for foundation models to acquire a deeper understanding of physical and household concepts. This learning necessitates a broader spectrum of information sources, such as diverse sensors (including smell, tactile, and thermal sensors), and more intricate data such as proprioception data from robot platforms with high degrees of freedom.
Current dexterous manipulators, e.g., Shadow Hand [230], are prohibitively expensive and prone to frequent breakdowns, hence they are predominantly experimented with in simulation. Moreover, tactile sensors are still limited in their application, often confined to the fingertips, as in [231], or offer only low resolution, as observed in the robot-sweater [232]. Recent progress has started to explore full-hand, high-resolution tactile sensing for humanoid hands [233].
Furthermore, since the bulk of data-collection is still conducted through human demonstrations, platforms that facilitate more accurate and efficient data acquisition, such as ALOHA [234], AirExo [235] and Leap Hands [210], are gaining popularity. Therefore, we posit that significant contributions are yet to be made—not only in terms of software innovations, but also in hardware. These advancements are crucial for providing richer data-collection and, thus, expanding the conceptual space of robotics foundation models.
@@
The robotics community needs to encourage standardized and reproducible research practices to ensure that published findings can be validated and compared by others. To enable reproducibility at scale, we need to bridge the gap between simulated environments and real-world hardware and improve the transferability of ML models. Homerobot [100] is a promising step towards enabling both simulation and hardware platforms for open vocabulary pick-and-place tasks. However, we need to establish standardized task definitions and affordances to handle different robot morphologies, enabling more efficient model development.
Simulation or Real-world Data Collection
The path to robotic foundation models could be diverse. One method to achieve this is through large-scale real-world collected data  [67, 253]. Another approach, however, is though simulation data and reduce the sim-to-real gap  [254]. Simulation has the advantage of providing theoretically unlimited amount of data, training robot learning policies without worrying about safety concerns, and potential efficient training via parallelism. The large-scale training on simulation has also shown good generalization abilities in robot policy learning [255] and perception models [256]. However, building the twin version of the real world also faces various of challenges. We look forward to seeing the potential of both these two approaches to robotic foundation models.
Deployability to Industrial Settings
Discussions in previous subsections have primarily focused on academic settings, but foundation models for robotics have significant potential to transform unstructured industrial environments such as construction, oil and gas, solar farms, mines, and agriculture (see Fig. 8). These environments are dynamic, large in scale, and can lack prior maps or external communication (e.g., can be GPS-denied). Additionally, hazardous materials, physical constraints like background noise, and limited visibility (e.g., masks and helmets) pose significant human-robot interaction challenges that bring to surface novel problems that must be addressed in such settings.
Figure 8: Industrial environments present unique challenges like dynamic obstacles, unstructured terrains, and safety risks. Foundation models in robotics offer the potential to navigate, inspect, and optimize operations in such demanding settings, improving safety and efficiency while enabling real-time decision-making and interaction with existing machinery.
Foundation models could revolutionize industrial operations by enabling robots navigate and inspect demanding outdoor settings, while supporting applications including automated data analysis [257], natural language processing for actionable insights [258], automated documentation, job site activity recognition [259], field operations optimization [260], and safety violation detection [261]. For instance, ConceptFusion [262] and Open-Fusion [263] enable open-set multimodal 3D mapping capabilities, potentially useful for dynamically-changing industrial settings where predefined object categories are unavailable. Such advancements can enable more flexible and adaptive robot performance in unstructured settings.
However, deploying these models in industrial contexts presents unique challenges. Ensuring safety and risk awareness to meet stringent industrial safety standards is crucial. Industrial robots often must adapt to dynamic obstacles and situations, varying surface conditions, uneven slopes, weather conditions, while maintaining awareness of potential hazards, including heavy machinery, all of which must be managed by the foundation model. For example, in construction environments, robots must navigate ever-changing sites due to ongoing project development, all while making safety-critical decisions [264]. In mining, such foundation models must handle adverse conditions including low visibility, variable terrains, high dust levels, and unstable surfaces [265]. The variability of these real-world environments means that foundation models deployed to such settings must operate in out-of-distribution settings, increasing the risk of unsafe decisions and erroneous outputs. ConceptGraphs [194], for example, can provide a framework to represent aspects of such environments using 3D scene graphs, while methods inspired by PlayFusion [266], which uses language annotation of unstructured / unguided robotics interaction data, bear potential to support robots in acquiring the necessary skills to interact with machinery.
Deployment also requires edge computing due to potential network limitations and latency while on-site. These mobile robots, are also constrained by computational resources, operating and collecting data for real-time monitoring. Techniques such as model quantization [267], distillation [268], and teacher-student based frameworks [269, 270] can help address these constraints, though balancing performance with model size remains an ongoing challenge.
@@ 6.2 Summary
Meta Analysis Tables
Below is a detailed analysis answering the questions raised in section 5.2. We list all the tables by classifying the topics as manipulation in Table 2, dexterous manipulation 3, mobile manipulation 4, locomotion 5, navigation 6 and multi-task learning 7.
@@
Due to the rapidly changing nature of the field, we checkpointed this version of literature review on September 1st 2024, and might have missed some relevant work. In addition, due to the rich body of literature and the extensiveness of this survey, there may be inaccuracies or mistakes in the paper. We welcome readers to send pull requests to our GitHub repository (inside https://robotics-fm-survey.github.io/) so we may continue to update our references, correct the mistakes and inaccuracies, as well as updating the entries of the meta studies in the paper. Please refer to the contribution guide in the GitHub repository.
@@ Acknowledgments
We would like to thank Vincent Vanhoucke for feedbacks on a draft of this survey paper. In addition, we would like to thank Yu Quan Chong and Kedi Xu for insightful discussions about the papers list.
Table 2: Tabletop Manipulation
Title Datasets & Simulation Real Robot Number of Tasks Base Model SR SR Descriptions Frequency RT-X Kitchen Manipulation [271], Cable Routing [272], NYU Door Opening [273], AUTOLab UR5 [274], Robot Play [275], Bridge [209], RL Bench (RT-1) [29] Google robot 160266 RT-2(PaLI) 30% 63% 92% Bridge, Small lab datasets, Google robot 3-10Hz RT-2 RL Bench (RT-1) [29] Language-Table [172] Google robot 480+ PaLI, PaLM-E 93% 62% Seen tasks, Unseen tasks 1-5Hz RT-1[29] RL Bench (RT-1) [29] Google robot 744 RT-1 97% 76% 83% 59% Seen Tasks, Unseen Tasks, With Distractors, Novel background 3 Hz GNFactor [99] RL Bench (RT-1) [29] XArm7 166 Stable Diffusion, CLIP 32% 28% Multi-task test, Unseen tasks Unspec. MOO [157] Self-created robotic data Google robot 106 Owl-ViT, RT-1 92% 75% Seen Tasks, Unseen Tasks Unspec. PhysObjects [206] PhysObjects [206] Franka Panda 51 FlanT5-XXL, GPT-4 88% 88% PhysObjects Test Set, Real scenes open loop Matcha [276] CoppeliaSim [277] NiCOL 50 ViLD, text-davinci-003 91% 57% Distinct descriptions, Indistinct descriptions open loop Scalingup [142] Scalingup benchmark [142] UR5 arm 18 GPT-3 79% Mean Success 35 Hz F3RM [98] PyBullet [278] and Self-created robotic data Franka Panda 18 CLIP 78% 62% Grasp-Place, Object Manip. Unspec. VIMA [171] VIMABench [171] None 17 pre-trained T5 tokenizer 81% 81% 79% 49% L1, L2, L3, L4 Unspec. Instruct2Act [279] VIMABench [171] None 17 CLIP, SAM, text-davinci-003, LLaMA 84% Mean Success open loop VoxPoser [129] Saipen [280] Where2act [281] Franka Emika 13 GPT-4, SAM, Owl-ViT 88% 70% 67% Static environments, With disturbances, Unseen semarios 5HZ Text2Motion [120] TableEnv [120] Franka Panda 6 text-davinci-003 82% Mean Success open loop GenSim [123] PyBullet [278] and Self-created robotic data XArm7 100+ GPT-4 53.3% 68.8% 50 Tasks 70 Tasks Unspec.
Table 3: Dexterous Manipulation
Title Datasets & Simulation Real Robot Number of Tasks Base Model SR SR Descriptions Frequency RoboCat [92] RGB-Stacking Benchmark [282] Sawyer 7-DoF, Panda 7-DoF, KUKA 14-DoF 253 RoboCat 82%, 74%, 86% Sawyer in sim, Panda in sim, KUKA in real 10 Hz 20 Hz
Table 4: Mobile Manipulation
Title Datasets & Simulation Real Robot Number of Tasks Base Model SR SR Descriptions Frequency LLaRP [283] Language Rearrangement [283] None 1000 LLaMA-7B 42% 0% 8% 39% Total, Multiple objects, Spatial Reasoning, Conditional instuct. open loop Code-As-Policies [122] Customized RoboCodeGen [122], HumanEval [284] Eveyday Robots, UR5 Robot arm 214 Codex, GPT-3 95% 96% RoboCodeGen, HumanEval open loop InnerMonlogue [154] Ravens [285]; UR5e robot, Google robot 130 ViLD GPT-3.5 51% 50% 90% 60% Seen Ravens tasks, Unseen Ravens tasks, Real robot arm, Google robot Unspec. LIV [286] MetaWorld [287] FrankaKitchen [183] Franka panda 114 LIV 30% 30% 45% FrankaKitchen, MetaWorld, Real robots 15hz SayCan [26] proprietary simulator and self-created robotic data Google robot 101 PaLM, Flan 74% Mean Success open loop PaLM-E [33] Lang-table [172] Google robot, xArm6 100 PaLM-E 83% 76% 52% 91% Grasping, Stacking, Lang-table tasks, Google robot 1-5Hz TidyBot [288] TidyBot [288] TidyBot 96 ViLD, CLIP , text-davinci003 91% 83% Sim, Real open loop LLM-Grop [289] Gazebo [290] Segway base + UR5e arm 8 GPT-3 4.08 Average human rating(1-5) open loop LLM+P [153] International Planning Competition [125] None 7 GPT-4 20% 90% 0% 95% 85% 20% 10% Barman, BLOCKSWORLD, FLOORTILE, GRIPPERS, STORAGE, TERMES, TYREWORLD open loop HomeRobot [100] Habitat [291] Hello-Robot Stretch 8 - 20% RL baseline closed loop
Table 5: Locomotion
Title Datasets & Simulation Real Robot Number of Tasks Base Model SR SR Descriptions Frequency SayTap [28] IsaacGym [62] Unitree A1 30 GPT-4 97% Mean Success openloop Prompt2Walk [130] Mujoco [61] None 1 GPT-4 80% Mean Success 10 Hz TokenHumanoid2024 [292] IsaacGym [62], Agility Robotics’ simulator Agility Robotics Humanoid 15 GPT-4 Tracking error & Prediction error openloop
Table 6: Navigation
Title Datasets & Simulation Real Robot Number of Tasks Base Model SR SR Descriptions Frequency LM-Nav [96] Self-created datasets Clearpath Jackal UGV 20 GPT-3 80% Mean Success open loop CLIP-Fields [94] Habitat-Matterport 3D Semantic [291] Stretch Robot 14 CLIP Detic 79% Mean Success open loop GNM [191] GNM datasets [191] LoCoBot Vizbot DJI Tello Jackal UGV 3 GNM 96% Mean Success Unspec.
Table 7: Multi-Tasks
Title Datasets & Simulation Real Robot Number of Tasks Base Model SR SR Descriptions Frequency Gato [32] Meta-World [287], Sokoban [293], BabyAI [294], Procgen Benchmark [295], Arcade Learning Environment [296], DM Control Suite [297] Sawyer arm 604 GPT-4 97% 68% 80% 68% super-human 64% Meta-World, Sokoban, BabyAI, Procgen, ALE Atari, DM Control Suite 20 Hz Grounded Decoding [298] 2D Maze [299], Ravens [285] None 124 GPT-3.5 PaLM 71% 46% 95% 51% Ravens Seen Tasks, Ravens Seen Tasks, 2D Maze, Mobile Manipulation open loop Eureka [78] Issac Gym [62], Bidexterous Manipulation [300] None 29 GPT-4 55% 3.7 1 Bi-Dextrous, Isaac Tasks – (Human Score) open loop Lang2Reward [134] MuJoCo MPC [301] Google robot 17 GPT-4 95% 80% Quadruped, Dextrous manipulator open loop VC-1 [187] CortexBench [187] None 17 VC-1 59% 89% 67% 72% 60% 70% 63% Adroit, Meta-World, DMControl, Trifinger, ObjectNav, ImageNav, Mobile Pick open loop